OpenStudy (anonymous):

Solve -2x-1=(3x+5)/2 for x. Show all work.

myininaya (myininaya):

no me... multiply by 2 on both sides -4x-2=3x+5 add 4x on both sides -2=7x+5 subtract 5 on both sides -9=7x divide by 7 on both sides -9/7=x

myininaya (myininaya):

winner!

OpenStudy (anonymous):

\[-2x-1=\frac{3x+5}{2}\] \[2(-2x-1)=3x+5\] \[-4x-2=-3x+5\] \[-2-5=-3x+4x \] \[-7=x\]

OpenStudy (anonymous):

oops i lose and i am wrong!!!!

OpenStudy (anonymous):

replaced 3x by -3x

myininaya (myininaya):

haha

OpenStudy (anonymous):

myininaya we will have another race when you write the latex!

myininaya (myininaya):

lol